A model of a helicopter rotor has four blades, each 3.40 m long from the central shaft to the blade tip. The model is rotated in a wind tunnel at 550 rev/min.


(a) What is the linear speed of the blade tip, in m/s?


(b) What is the radial acceleration of the blade tip expressed as a multiple of the acceleration?


1
Expert's answer
2021-09-21T14:18:10-0400

a) The linear speed in circular motion is given as follows:


"v = 2\pi fr"

where "f = 550rev/min = \dfrac{55}{6}rev/min" is the frequency, and "r=3.4m" is the radius. Thus, obtain:


"v = 2\pi\cdot \dfrac{55}{6}\cdot 3.4 \approx 196m/s"

b) The radial acceleration is given as follows:


"a = \dfrac{v^2}{r} = \dfrac{196^2}{3.4} \approx 57.6m/s^2"

Answer. a) 196 m/s, b) 57.6 m/s^2.